# PoSh: Using Scene Graphs to Guide LLMs-as-a-Judge for Detailed Image Descriptions

### v2

@@ 1 Introduction
A picture is worth a thousand words – can vision-language models (VLMs) capture all of them? VLMs have saturated traditional image understanding benchmarks from short captioning to question answering (li2025benchmark). New, more challenging tasks are needed to measure VLM progress. Detailed image description is of particular interest as it requires comprehensive understanding – e.g., in Fig. 1, a VLM must correctly specify who is pouring the water. This deep perception is a better proxy for the demands of the real world, where diverse user queries may not be reflected in VQA benchmarks (chenwe). Moreover, it enables meaningful applications such as image assistive (“alt”) text generation that could greatly expand accessibility online (mack2021we).
However, making progress on detailed description requires cheap, reliable methods for scoring models. Human evaluation is costly, involving the painstaking comparison of long texts. Even so, there is often no substitute as most metrics were designed for short texts and older models (berger2024surveying). Moreover, while metrics that produce a single coarse score of overall quality allow for the ranking of models, they offer little insight into the granular issues driving performance. Granular issues include mistakes in each generation, like the positions of the people in Fig. 1, and omissions in each reference, like the details of the bird’s beak in Fig. 3. Automatically localizing such errors is critical as long generations with similar coarse scores may differ in multiple dimensions of interest (e.g., facial features, body orientations, etc.). Otherwise, prompt and/or model iteration necessitates expensive manual inspection to understand which description aspects need improvement.
@@
In this work, we propose PoSh  [^1] PoSh (PrOofing Scene grapHs) can judge if your detailed descriptions are what you (really really) want., a metric for evaluating detailed descriptions that addresses these challenges. PoSh extracts scene graphs from a generated description and its reference to use as structured rubrics for an LLM to granularly identify mistakes and omissions (see Fig. 2), pinpointing the textual spans containing errors like attribute/relation mis-attachment. Then, it aggregates these localized errors into coarse scores for mistakes, omissions and overall quality. Thus, PoSh weds the strengths of structured methods like scene graphs (anderson2016spice), which reduce descriptions to their consequential visual components, with the strengths of LLMs/VLMs-as-a-Judge (zheng2023judging), which flexibly compare these visual components against diverse surface realizations.
@@
We validate PoSh against the human judgments in DOCENT. We show that PoSh recovers human description rankings more often (+$3$ percentage points) and achieves stronger correlations with human-derived scores (+$0.05$ Spearman $\rho$) than existing overlap and open-weight alternatives (e.g. SPICE, CAPTURE, LLaVa-Critic), even surpassing GPT4o-as-a-Judge. Moreover, using judgments in CapArena (cheng2025caparena), we show this strength is robust to image type. Then, given its calibration, we experiment with using PoSh as a reward function for describing the images in DOCENT and find that this yields meaningfully better descriptions than supervised fine-tuning (SFT).
Finally, using PoSh, we characterize the performance of open and closed models in describing the artwork in DOCENT, establishing a difficult new task. In so doing, we extend detailed description to a technically challenging and socially impactful domain: assistive text generation for artwork, whose visual complexity and diversity stress VLMs (bengamra2024comprehensive) (see Fig. 1).
@@ 2 Related Work
Image description is under-specified – the correct way to describe an image is often task-specific. This is especially true for assistive text which has context-dependent requirements (kreiss2022context). Moreover, in such sensitive applications, correlated failures between reference-free metrics and VLMs relying on similar components could prove dangerous to end users (deutsch2022limitations). Thus, our focus is reference-based evaluation. Traditional metrics were not designed to evaluate long text and can involve truncation due to limited context length (e.g. CLIPScore) (papineni2002bleu; lin2004rouge; banerjee2005meteor; vedantam2015cider; see2017get; hessel2021clipscore; sarto2023positive). Recent work has explored LLMs/VLMs-as-Judges though this requires potentially expensive API calls and offers limited replicability (chan2023clair; cheng2025caparena). Even when replicable, they do not provide interpretable, grounded granular scores (xiong2024llavacritic).
While prior metrics like SPICE and CAPTURE leverage scene graphs, they forgo their rich structure by ignoring object attachment (anderson2016spice; dong2024benchmarking). This favors generations with misattributed details (as in Fig. 1). In summarization, scialom2021questeval use question generation and answering (QA) to compare a summary and its source. In text-to-image generation, chodavidsonian use GPT4 to extract and verify a scene graph from a visual prompt. PoSh extends these approaches to detailed description evaluation that is replicable and interpretable. With small models, it extracts scene graphs to use as structured rubrics for guiding an open-weight LLM-as-a-Judge.
@@
Evaluating such a metric requires human judgments of model generations. Though there are many detailed image description benchmarks (urbanek2024picture; onoe2024docci; garg-etal-2024-imageinwords; lu2025benchmarking; ye2025painting), most release no such judgments. One notable exception is CapArena which contains coarse rankings of descriptions for web imagery (cheng2025caparena). In contrast, our new dataset, DOCENT contains both granular and coarse judgments, enabling the evaluation of fine-grained metrics like PoSh. Moreover, it expands detailed description to artwork whose scene dynamics and expert-written references are considerably more complex (see Table 1).
@@ 3 PoSh: A New Metric for Detailed Image Description
As in SPICE (anderson2016spice), given a description $d$, a scene graph $G(d)$ is a structured representation of $d$. Specifically, $G(d)=\left<O(d),E(d),K(d)\right$> where $O(d)\subseteq C$ is a set of objects, $E(d)\subseteq O(d)\times A$ is a set of attributes associated with each object and $K(d)\subseteq O(d)\times R\times O(d)$ are a set of relation edges between objects. $C$, AA and $R$ are open-world sets of all possible object, attribute and relation classes.
Given a generation gen with its reference ref we extract sentence-level scene graphs $G_{i}(\textit{gen}),G_{j}(\textit{ref})$ for each using off-the-shelf dependency parsing and combine them via coreference resolution (spacy; martinelli-etal-2024-maverick). This produces scene graphs with full coverage of each text where each component is localized to text spans, allowing for grounded, interpretable scoring. We provide pseudocode for this extraction in Section A.1.2.
@@
We note this is a natural place to introduce tunable weights (as in dong2024benchmarking) to adapt PoSh to particular datasets. As we aim to demonstrate robustness, we leave these terms unweighted.
@@ 4 DOCENT: A New Benchmark for Detailed Description of Art
DOCENT is a benchmark for evaluating detailed description metrics and detailed descriptions themselves. It consists of $1,750$ works of art with expert-written references from the Open Data Program at the U.S. National Gallery of Art (NGA) [^4] https://www.nga.gov/open-access-images/open-data.html. For $100$ of these images, we produce four generations from current small and frontier VLMs and collect $300$ granular (for $75$ images) and $600$ coarse judgments from annotators knowledgeable in art of mistakes and omissions [^5] We forgo fluency as recommended by kasai-etal-2022-transparent. On average, coarse judgments took $5$ minutes and granular judgments took $18$ minutes (six annotation days). This highlights both the cost of manual evaluation and the need for metrics that are reliable proxies.
@@
The accompanying references are detailed descriptions whose purpose is accessibility – as such, they follow guidelines666www.nga.gov/visit/accessibility/collection-image-descriptions that include tips for describing color (e.g., “color can be likened to temperature”) and handling ambiguity (e.g, “describe what makes something ambiguous”). These context informed requirements highlight the need for reference based metrics (kreiss2022context).
Compared to existing detailed image description benchmarks, DOCENT contains considerably more visual complexity (see Table 1). On average, its images contain $16\%$ more objects and nearly twice as many people [^7] As measured by OneFormer jain2023oneformer who require description of their orientation, features, clothing, etc. Consequently, the average length and scene graph size of its reference descriptions are nearly double.
@@
We generate detailed descriptions for $100$ images in DOCENT from four current VLMs that span transparency and model size (from open data/open weight to frontier models): LLaVA-1.6-7B (liu2024improved), Molmo-D-7B (deitke2024molmo), GPT4o and Claude Sonnet 3.5. A metric that discriminates among these generations similarly to their human judgments could gauge progress in detailed image description in small and large VLMs over time. Additional details (prompts, date of API access) can be found in Section A.2.1.
@@
Half of our annotators identify mistakes and omissions in our model generations. For each image, an annotator is shown its reference and then its four model generations in random order. First, they look at the image, read the reference and then the current generation. Next, by selecting narrow text spans, annotators first identify mistakes in the generation (i.e. precision errors) and then omissions in the reference that are not in the generation (i.e. recall errors). When identifying omissions, as in kasai-etal-2022-transparent, we ask annotators to mentally correct narrow mistakes in the generation first to avoid double-penalizing a model for both incorrect specificity and lack of specificity. For example, a generation that describes a woman as a man is an error in precision but not in recall. We include our task instructions and interface [^9] Hosted on Label Studio (https://labelstud.io) screenshots in Figs. 4 and 6.
@@
The other half of our annotators provide coarse judgments of our model generations. For a given image, an annotator is shown its reference and two generations (#1 and #2) in random order and asked to rank the generations in terms of mistakes (i.e. precision), omissions (i.e. recall) and overall quality. These pairwise judgments avoid some of the inter-annotator inconsistency of Likert ratings, especially for long text (novikova2018rankme).
@@
For our granular judgments, as recommended by hripcsak2005agreement for span annotation tasks where the boundaries of negative examples (i.e. non-errors) are ill-defined, we measure agreement using the relaxed F1 (matching spans that contain 50% overlapping tokens). Under this measure, our student annotators exhibit strong agreement among themselves and with our expert.
Our coarse judgments exhibit moderate inter-annotator agreement, with Krippendorf’s $\alpha=0.509$, $0.409$ and $0.459$ for mistakes, omissions and overall quality (landis1977measurement). This level of agreement is unsurprising for coarse detailed description evaluation – judgment requires weighing the relative importance of each text’s granular errors and is consequently more subjective. Nevertheless, our student annotators exhibit moderate to strong correlations with our expert, with significant Pearson $\rho$ values of $0.727$, $0.501$ and $0.492$ for mistakes, omissions and overall quality respectively.
@@ 5 Experiments
We extract sentence-level scene graphs using en_core_web_trf from spacy, a transformer trained to perform dependency parsing. To merge objects across these scene graphs while preserving attribute and relation attachments, we use maverick-mes-ontonotes from martinelli-etal-2024-maverick to perform co-reference resolution. Our QA scorer $\Psi$ is qwen-3-14b (yang2025qwen3). We template evaluation questions for each scene graph component (as in Fig. 2), re-write candidate identifiers using $\Psi$ to improve fluency and then prompt $\Psi$ to answer each templated presence question by predicting a number between $1$ and $5$. We extract scores by taking the weighted average over the token logits for each number as in liu2023g. When determining object presence, we use a threshold of $2$, determined through tuning on a small hand-annotated validation set. We provide further implementation details and all prompts used in Section A.1.
@@
DOCENT is our new detailed description benchmark containing judgments from knowledgeable human annotators: granular mistake and omission spans for $300$ individual generations and coarse scaled rankings of mistakes, omissions and overall quality of $600$ paired generations. We evaluate granular metrics on this benchmark using macro F1 where we credit/penalize a model for predicting each annotated/unannotated token. Our coarse judgments are in the form $(text_{1},text_{2},score)$ where score indicates how much better or worse $text_{1}$ is than $text_{2}$. We evaluate each coarse metric $m$ by calculating its 1) pairwise accuracy (whether it picks the better text or a tie, using a tie threshold inferred from the gold tie rate) and 2) Spearman rank $\rho$ and Kendall’s $\tau$ correlations between $m(text_{1})-m(text_{2})$ and scorescore, a common practice in machine translation metric evaluation (kocmi2021ship). More details can be found in Section A.3.1.
CapArena (cheng2025caparena) contains $3,361$ images and $10,348$ detailed descriptions generated from $14$ current VLMs. $5,599$ pairs of these generations receive coarse judgments from human raters of the better generation (or “tie”). We include CapArena, which contains diverse images drawn from the web, to validate metric robustness. However, we note the dramatic simplicity of its images and references compared to those in DOCENT (see Table 1). $64\%$ of its images [^10] As measured by OneFormer (jain2023oneformer) contain fewer than two objects and $95\%$ depict fewer than two people (compared to $27\%$ and $52\%$ in DOCENT). A metric is evaluated on CapArena at the caption-level (whether it picks the better text or a tie, using a tie threshold inferred from the gold tie rate) and at the model-level (Spearman’s rank and Kendall’s $\tau$ correlation between ELO rankings derived from metric predictions and gold judgments).
@@
Our work is the first to introduce both a metric and a benchmark for granular evaluation of detailed descriptions. As such, this limits our baselines to those able to predict localized mistakes and omissions (i.e., the spans where errors occur). We consider two embedding-based approaches, using Qwen/Qwen3-Embedding-8B from yang2025qwen3: 4GramEmbed, which embeds and compares $4$-grams from a generation and its reference, and SGEmbed, which embeds and compares components from the scene graphs of a generation and its reference. As these approaches (and PoSh) produce span scores, we report the maximum F1 scores for mistakes and omissions across all alerting thresholds. More details can be found in Section A.3.3.
@@
Though PoSh is a text-only reference-based metric, we select a representative set of reference-free (requiring only an image) and reference-based (requiring a gold standard) pointwise metrics (i.e. produce numerical scores) as our baselines. These include n-gram overlap metrics like BLEU (papineni2002bleu), ROUGE-L-Sum (see2017get), METEOR (banerjee2005meteor) and CIDER (vedantam2015cider) and model-based metrics like SPICE (anderson2016spice), CLIPScore (hessel2021clipscore) and CAPTURE (dong2024benchmarking). Additionally, we consider several LLMs/VLMs-as-a-Judge [^11] CLAIR/Faithscore were not included due to complications with their codebases (chan2023clair; jing2024faithscore). Due to cost (estimated at \$$1,000$), we only evaluate DCScore (ye2025painting) on DOCENT.: FLEUR (lee2024fleur), Prometheus (kim2023prometheus), LLaVA-Critic (xiong2024llavacritic), DCScore ye2025painting, Qwen-3 (yang2025qwen3) and GPT4o/GPT5 in three settings (reference-free with image, reference-based without image and reference-based with image). More details can be found in Section A.3.4.
@@
Finally, given the potential of a well-calibrated metric as a verifier in reinforcement learning (RL), we evaluate PoSh as a reward function. We train Qwen2.5-VL-7B on the $1,000$ images in DOCENT’s training set in two settings: 1) supervised fine-tuning (SFT), and 2) RL with DAPO (yu2025dapo) using PoSh. We collect coarse judgments (as in Section 4) for $40$ generation pairs from graduate students in NLP. More details can be found in Section A.3.5.

### v3

@@ 1 Introduction
A picture is worth a thousand words – can vision-language models (VLMs) capture all of them? VLMs have saturated traditional image understanding benchmarks from short captioning to question answering (Li et al., 2025). New, more challenging tasks are needed to measure VLM progress. Detailed image description is of particular interest as it requires comprehensive understanding – e.g., in Fig. 1, a VLM must correctly specify who is pouring the water. This deep perception is a better proxy for the demands of the real world, where diverse user queries may not be reflected in VQA benchmarks (Chen et al., 2024). Moreover, it enables meaningful applications such as image assistive (“alt”) text generation that could greatly expand accessibility online (Mack et al., 2021).
However, making progress on detailed description requires cheap, reliable methods for scoring models. Human evaluation is costly, involving the painstaking comparison of long texts. Even so, there is often no substitute as most metrics were designed for short texts and older models (Berger et al., 2024). Moreover, while metrics that produce a single coarse score of overall quality allow for the ranking of models, they offer little insight into the granular issues driving performance. Granular issues include mistakes in each generation, like the positions of the people in Fig. 1, and omissions in each reference, like the details of the bird’s beak in Fig. 3. Automatically localizing such errors is critical as long generations with similar coarse scores may differ in multiple dimensions of interest (e.g., facial features, body orientations, etc.). Otherwise, prompt and/or model iteration necessitates expensive manual inspection to understand which description aspects need improvement.
@@
In this work, we propose PoSh  [^1] PoSh (PrOofing Scene grapHs) can judge if your detailed descriptions are what you (really really) want., a metric for evaluating detailed descriptions that addresses these challenges. PoSh extracts scene graphs from a generated description and its reference to use as structured rubrics for an LLM to granularly identify mistakes and omissions (see Fig. 2), pinpointing the textual spans containing errors like attribute/relation mis-attachment. Then, it aggregates these localized errors into coarse scores for mistakes, omissions and overall quality. Thus, PoSh weds the strengths of structured methods like scene graphs (Anderson et al., 2016), which reduce descriptions to their consequential visual components, with the strengths of LLMs/VLMs-as-a-Judge (Zheng et al., 2023), which flexibly compare these visual components against diverse surface realizations.
@@
We validate PoSh against the human judgments in DOCENT. We show that PoSh recovers human description rankings more often (+$3$ percentage points) and achieves stronger correlations with human-derived scores (+$0.05$ Spearman $\rho$) than existing overlap and open-weight alternatives (e.g. SPICE, CAPTURE, LLaVa-Critic), even surpassing GPT4o-as-a-Judge. Moreover, using judgments in CapArena (Cheng et al., 2025), we show this strength is robust to image type. Then, given its calibration, we experiment with using PoSh as a reward function for describing the images in DOCENT and find that this yields meaningfully better descriptions than supervised fine-tuning (SFT).
Finally, using PoSh, we characterize the performance of open and closed models in describing the artwork in DOCENT, establishing a difficult new task. In so doing, we extend detailed description to a technically challenging and socially impactful domain: assistive text generation for artwork, whose visual complexity and diversity stress VLMs (Bengamra et al., 2024) (see Fig. 1).
@@ 2 Related Work
Image description is under-specified – the correct way to describe an image is often task-specific. This is especially true for assistive text which has context-dependent requirements (Kreiss et al., 2022). Moreover, in such sensitive applications, correlated failures between reference-free metrics and VLMs relying on similar components could prove dangerous to end users (Deutsch et al., 2022). Thus, our focus is reference-based evaluation. Traditional metrics were not designed to evaluate long text and can involve truncation due to limited context length (e.g. CLIPScore) (Papineni et al., 2002; Lin, 2004; Banerjee and Lavie, 2005; Vedantam et al., 2015; See et al., 2017; Hessel et al., 2021; Sarto et al., 2023). Recent work has explored LLMs/VLMs-as-Judges though this requires potentially expensive API calls and offers limited replicability (Chan et al., 2023; Cheng et al., 2025). Even when replicable, they do not provide interpretable, grounded granular scores (Xiong et al., 2025).
While prior metrics like SPICE and CAPTURE leverage scene graphs, they forgo their rich structure by ignoring object attachment (Anderson et al., 2016; Dong et al., 2024). This favors generations with misattributed details (as in Fig. 1). In summarization, Scialom et al. (2021) use question generation and answering (QA) to compare a summary and its source. In text-to-image generation, Cho et al. (2024) use GPT4 to extract and verify a scene graph from a visual prompt. PoSh extends these approaches to detailed description evaluation that is replicable and interpretable. With small models, it extracts scene graphs to use as structured rubrics for guiding an open-weight LLM-as-a-Judge.
@@
Evaluating such a metric requires human judgments of model generations. Though there are many detailed image description benchmarks (Urbanek et al., 2024; Onoe et al., 2024; Garg et al., 2024; Lu et al., 2025; Ye et al., 2025), most release no such judgments. One notable exception is CapArena which contains coarse rankings of descriptions for web imagery (Cheng et al., 2025). In contrast, our new dataset, DOCENT contains both granular and coarse judgments, enabling the evaluation of fine-grained metrics like PoSh. Moreover, it expands detailed description to artwork whose scene dynamics and expert-written references are considerably more complex (see Table 1).
@@ 3 PoSh: A New Metric for Detailed Image Description
As in SPICE (Anderson et al., 2016), given a description $d$, a scene graph $G(d)$ is a structured representation of $d$. Specifically, $G(d)=\left<O(d),E(d),K(d)\right$> where $O(d)\subseteq C$ is a set of objects, $E(d)\subseteq O(d)\times A$ is a set of attributes associated with each object and $K(d)\subseteq O(d)\times R\times O(d)$ are a set of relation edges between objects. $C$, AA and $R$ are open-world sets of all possible object, attribute and relation classes.
Given a generation gen with its reference ref we extract sentence-level scene graphs $G_{i}(\textit{gen}),G_{j}(\textit{ref})$ for each using off-the-shelf dependency parsing and combine them via coreference resolution (Honnibal et al., 2020; Martinelli et al., 2024). This produces scene graphs with full coverage of each text where each component is localized to text spans, allowing for grounded, interpretable scoring. We provide pseudocode for this extraction in Section A.1.2.
@@
We note this is a natural place to introduce tunable weights (as in Dong et al. (2024)) to adapt PoSh to particular datasets. As we aim to demonstrate robustness, we leave these terms unweighted.
@@ 4 DOCENT: A New Benchmark for Detailed Description of Art
DOCENT is a benchmark for evaluating detailed description metrics and detailed descriptions themselves. It consists of $1,750$ works of art with expert-written references from the Open Data Program at the U.S. National Gallery of Art (NGA) [^4] https://www.nga.gov/open-access-images/open-data.html. For $100$ of these images, we produce four generations from current small and frontier VLMs and collect $300$ granular (for $75$ images) and $600$ coarse judgments from annotators knowledgeable in art of mistakes and omissions [^5] We forgo fluency as recommended by Kasai et al. (2022). On average, coarse judgments took $5$ minutes and granular judgments took $18$ minutes (six annotation days). This highlights both the cost of manual evaluation and the need for metrics that are reliable proxies.
@@
The accompanying references are detailed descriptions whose purpose is accessibility – as such, they follow guidelines666www.nga.gov/visit/accessibility/collection-image-descriptions that include tips for describing color (e.g., “color can be likened to temperature”) and handling ambiguity (e.g, “describe what makes something ambiguous”). These context informed requirements highlight the need for reference based metrics (Kreiss et al., 2022).
Compared to existing detailed image description benchmarks, DOCENT contains considerably more visual complexity (see Table 1). On average, its images contain $16\%$ more objects and nearly twice as many people [^7] As measured by OneFormer Jain et al. (2023) who require description of their orientation, features, clothing, etc. Consequently, the average length and scene graph size of its reference descriptions are nearly double.
@@
We generate detailed descriptions for $100$ images in DOCENT from four current VLMs that span transparency and model size (from open data/open weight to frontier models): LLaVA-1.6-7B (Liu et al., 2024), Molmo-D-7B (Deitke et al., 2024), GPT4o and Claude Sonnet 3.5. A metric that discriminates among these generations similarly to their human judgments could gauge progress in detailed image description in small and large VLMs over time. Additional details (prompts, date of API access) can be found in Section A.2.1.
@@
Half of our annotators identify mistakes and omissions in our model generations. For each image, an annotator is shown its reference and then its four model generations in random order. First, they look at the image, read the reference and then the current generation. Next, by selecting narrow text spans, annotators first identify mistakes in the generation (i.e. precision errors) and then omissions in the reference that are not in the generation (i.e. recall errors). When identifying omissions, as in Kasai et al. (2022), we ask annotators to mentally correct narrow mistakes in the generation first to avoid double-penalizing a model for both incorrect specificity and lack of specificity. For example, a generation that describes a woman as a man is an error in precision but not in recall. We include our task instructions and interface [^9] Hosted on Label Studio (https://labelstud.io) screenshots in Figs. 4 and 6.
@@
The other half of our annotators provide coarse judgments of our model generations. For a given image, an annotator is shown its reference and two generations (#1 and #2) in random order and asked to rank the generations in terms of mistakes (i.e. precision), omissions (i.e. recall) and overall quality. These pairwise judgments avoid some of the inter-annotator inconsistency of Likert ratings, especially for long text (Novikova et al., 2018).
@@
For our granular judgments, as recommended by Hripcsak and Rothschild (2005) for span annotation tasks where the boundaries of negative examples (i.e. non-errors) are ill-defined, we measure agreement using the relaxed F1 (matching spans that contain 50% overlapping tokens). Under this measure, our student annotators exhibit strong agreement among themselves and with our expert.
Our coarse judgments exhibit moderate inter-annotator agreement, with Krippendorf’s $\alpha=0.509$, $0.409$ and $0.459$ for mistakes, omissions and overall quality (Landis and Koch, 1977). This level of agreement is unsurprising for coarse detailed description evaluation – judgment requires weighing the relative importance of each text’s granular errors and is consequently more subjective. Nevertheless, our student annotators exhibit moderate to strong correlations with our expert, with significant Pearson $\rho$ values of $0.727$, $0.501$ and $0.492$ for mistakes, omissions and overall quality respectively.
@@ 5 Experiments
We extract sentence-level scene graphs using en_core_web_trf from Honnibal et al. (2020), a transformer trained to perform dependency parsing. To merge objects across these scene graphs while preserving attribute and relation attachments, we use maverick-mes-ontonotes from Martinelli et al. (2024) to perform co-reference resolution. Our QA scorer $\Psi$ is qwen-3-14b (Yang et al., 2025). We template evaluation questions for each scene graph component (as in Fig. 2), re-write candidate identifiers using $\Psi$ to improve fluency and then prompt $\Psi$ to answer each templated presence question by predicting a number between $1$ and $5$. We extract scores by taking the weighted average over the token logits for each number as in Liu et al. (2023). When determining object presence, we use a threshold of $2$, determined through tuning on a small hand-annotated validation set. We provide further implementation details and all prompts used in Section A.1.
@@
DOCENT is our new detailed description benchmark containing judgments from knowledgeable human annotators: granular mistake and omission spans for $300$ individual generations and coarse scaled rankings of mistakes, omissions and overall quality of $600$ paired generations. We evaluate granular metrics on this benchmark using macro F1 where we credit/penalize a model for predicting each annotated/unannotated token. Our coarse judgments are in the form $(text_{1},text_{2},score)$ where score indicates how much better or worse $text_{1}$ is than $text_{2}$. We evaluate each coarse metric $m$ by calculating its 1) pairwise accuracy (whether it picks the better text or a tie, using a tie threshold inferred from the gold tie rate) and 2) Spearman rank $\rho$ and Kendall’s $\tau$ correlations between $m(text_{1})-m(text_{2})$ and scorescore, a common practice in machine translation metric evaluation (Kocmi et al., 2021). More details can be found in Section A.3.1.
CapArena (Cheng et al., 2025) contains $3,361$ images and $10,348$ detailed descriptions generated from $14$ current VLMs. $5,599$ pairs of these generations receive coarse judgments from human raters of the better generation (or “tie”). We include CapArena, which contains diverse images drawn from the web, to validate metric robustness. However, we note the dramatic simplicity of its images and references compared to those in DOCENT (see Table 1). $64\%$ of its images [^10] As measured by OneFormer (Jain et al., 2023) contain fewer than two objects and $95\%$ depict fewer than two people (compared to $27\%$ and $52\%$ in DOCENT). A metric is evaluated on CapArena at the caption-level (whether it picks the better text or a tie, using a tie threshold inferred from the gold tie rate) and at the model-level (Spearman’s rank and Kendall’s $\tau$ correlation between ELO rankings derived from metric predictions and gold judgments).
@@
Our work is the first to introduce both a metric and a benchmark for granular evaluation of detailed descriptions. As such, this limits our baselines to those able to predict localized mistakes and omissions (i.e., the spans where errors occur). We consider two embedding-based approaches, using Qwen/Qwen3-Embedding-8B from Yang et al. (2025): 4GramEmbed, which embeds and compares $4$-grams from a generation and its reference, and SGEmbed, which embeds and compares components from the scene graphs of a generation and its reference. As these approaches (and PoSh) produce span scores, we report the maximum F1 scores for mistakes and omissions across all alerting thresholds. More details can be found in Section A.3.3.
@@
Though PoSh is a text-only reference-based metric, we select a representative set of reference-free (requiring only an image) and reference-based (requiring a gold standard) pointwise metrics (i.e. produce numerical scores) as our baselines. These include n-gram overlap metrics like BLEU (Papineni et al., 2002), ROUGE-L-Sum (See et al., 2017), METEOR (Banerjee and Lavie, 2005) and CIDER (Vedantam et al., 2015) and model-based metrics like SPICE (Anderson et al., 2016), CLIPScore (Hessel et al., 2021) and CAPTURE (Dong et al., 2024). Additionally, we consider several LLMs/VLMs-as-a-Judge [^11] CLAIR/Faithscore were not included due to complications with their codebases (Chan et al., 2023; Jing et al., 2024). Due to cost (estimated at \$$1,000$), we only evaluate DCScore (Ye et al., 2025) on DOCENT.: FLEUR (Lee et al., 2024), Prometheus (Kim et al., 2023), LLaVA-Critic (Xiong et al., 2025), DCScore Ye et al. (2025), Qwen-3 (Yang et al., 2025) and GPT4o/GPT5 in three settings (reference-free with image, reference-based without image and reference-based with image). More details can be found in Section A.3.4.
@@
Finally, given the potential of a well-calibrated metric as a verifier in reinforcement learning (RL), we evaluate PoSh as a reward function. We train Qwen2.5-VL-7B on the $1,000$ images in DOCENT’s training set in two settings: 1) supervised fine-tuning (SFT), and 2) RL with DAPO (Yu et al., 2025) using PoSh. We collect coarse judgments (as in Section 4) for $40$ generation pairs from graduate students in NLP. More details can be found in Section A.3.5.